OpenStudy (anonymous):

Solve for x. 3(3x - 1) + 2(3 - x) = 0

OpenStudy (anonymous):

x=3/7

OpenStudy (anonymous):

no... it's -3/7

OpenStudy (johnweldon1993):

distribute it 3(3x - 1) + 2(3 - x) = 0 (9x - 3) + (6 -2x) = 0 subtract (6-2x) from both sides 9x - 3 = -6 +2x subtract 2x from both sides 7x - 3 = -6 add 3 to both sides 7x = -3 divide both sides by 7 x = -3/7